OpenStudy (babyslapmafro):

Please help me with the following problem. A single degree of freedom mechanical system has a response with time that is harmonic. Based on your experimental data, you determine that the response has an amplitude of 1 cm and a frequency of 10 Hz. Determine the natural period, maximum velocity, and maximum acceleration.

OpenStudy (babyslapmafro):

I've only been able to solve for the natural period. \[\tau=\frac{ 1 }{ f }=\frac{ 1 }{ 10Hz }=0.1s\]

OpenStudy (phi):

model the position using \[ x= A \sin 2 \pi f \ t \] with A= 1 cm and f= 10 Hz, you have \[x = \sin 20 \pi t\]

OpenStudy (phi):

as your "system" oscillates back and forth (with a center at x=0), it will be slowest at the "edges" (think of a swing going back and forth) it will be fastest as it zooms by x=0 do you know calculus? if so , you can find \[ v= 20 \pi \cos(20 \pi t) \] and at t=0 or 0.05 or 0.1 you will have a peak velocity \( \pm 20 \pi\ cm/s\)